OpenStudy (madeleine2):

If the length of one leg of a right triangle is 16 and the length of the other leg is odd, what are the lengths of all the sides of the triangle? Other leg = Hypotenuse =

satellite73 (satellite73):

try solving \[2mn=16\] for whole numbers \(m\) and \(n\)

satellite73 (satellite73):

did you find anything yet?

OpenStudy (madeleine2):

im not sure what to do @satellite73

satellite73 (satellite73):

divide both sides by 2 would be a good first step

OpenStudy (madeleine2):

mm=8

OpenStudy (madeleine2):

then what @satellite73

satellite73 (satellite73):

think of a way to factor 8 as the product of two whole numbers, one of which is even, and one if which is odd

OpenStudy (madeleine2):

idk how to do that @satellite73

satellite73 (satellite73):

think is all what two whole numbers multiply to get 8?

satellite73 (satellite73):

did you think of two number \(m\) and \(n\) with \(m\times n=8\)?

OpenStudy (madeleine2):

4*2=8 or 1*8=8

OpenStudy (madeleine2):

@satellite73

satellite73 (satellite73):

yes good

satellite73 (satellite73):

now only one of those pairs has an even number and an odd number which one?

OpenStudy (madeleine2):

thanks @satellite73

satellite73 (satellite73):

wait we are not done yet at all

satellite73 (satellite73):

we are only gettting started